OpenStudy (anonymous):

need help solving \[(\sqrt{3}-2)(4\sqrt{12}+5)\]

OpenStudy (anonymous):

Simplifying \[\sqrt{12}\] gets you \[\sqrt{4\times3}\]

OpenStudy (anonymous):

Go from there

OpenStudy (anonymous):

\[(\sqrt{3}-)(8\sqrt{3}+5)\]now i expand the equation

OpenStudy (anonymous):

\[14-11\sqrt{3}\]

OpenStudy (anonymous):

\[8\sqrt{3}-8\sqrt{3}-16\sqrt{3}-10\]

OpenStudy (anonymous):

there is no 5root3?

OpenStudy (anonymous):

Oh sorry

OpenStudy (anonymous):

\[8\sqrt{3}+5\sqrt{3}-16\sqrt{3}-10\]

OpenStudy (anonymous):

So, whats your result?

OpenStudy (anonymous):

okay from that i am getting 14-11root3

OpenStudy (anonymous):

\[(8+5-16)\sqrt{3} -10\]

OpenStudy (anonymous):

oooh ok -3root3-10

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

No problem (: